40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Aug 22 02:40:12 2013, Jenne, Update, LSC, DRMI Locked for 1+ minute!!!!!! DRMI_First_Lock_1plusMinute_21Aug2013.png
    Reply  Thu Aug 22 07:57:57 2013, Lisa, Update, LSC, DRMI Locked for 1+ minute!!!!!! 
       Reply  Thu Aug 22 13:03:40 2013, Jenne, Update, LSC, DRMI Locked for 1+ minute!!!!!! 
    Reply  Thu Aug 22 10:20:32 2013, kiwamu, Update, LSC, DRMI Locked for 1+ minute!!!!!! 
    Reply  Thu Aug 22 13:20:54 2013, Koji, Update, LSC, DRMI Locked for 1+ minute!!!!!! 
    Reply  Sat Aug 24 00:11:07 2013, Koji, Update, LSC, DRMI Locked with improved lock streatches Screenshot-Untitled_Window.png
       Reply  Sun Aug 25 06:07:11 2013, rana, Update, LSC, DRMI Locked with improved lock streatches 
Message ID: 9061     Entry time: Sun Aug 25 06:07:11 2013     In reply to: 9060
Author: rana 
Type: Update 
Category: LSC 
Subject: DRMI Locked with improved lock streatches 

 We're ready for using the auto configure.

We can put our scripts for the MICH, PRMI, and DRMI into the IFO CONFIGURE screens for now and then it should be easy to get them into the Guardian once Jamie has the bugs worked out.

This screen can also be used to setup and start the dither alignment for each configuration (once we have one working for DRMI / SRM).

Also, now that the notches/bandstop filters for the violin modes have been move from the SUS into the LSC, we should fix the triggering to engage them a few seconds after the boosts.

ELOG V3.1.3-